# Ankita Das and Yogesh Simmhan

### v1

@@ Abstract
Edge devices have typically been used for DNN inferencing. The increase in the compute power of accelerated edges is leading to their use in DNN training also. As privacy becomes a concern on multi-tenant edge devices, Docker containers provide a lightweight virtualization mechanism to sandbox models. But their overheads for edge devices are not yet explored. In this work, we study the impact of containerized DNN inference and training workloads on an NVIDIA AGX Orin edge device and contrast it against bare metal execution on running time, CPU, GPU and memory utilization, and energy consumption. Our analysis provides several interesting insights on these overheads.
@@ 1 Introduction
    We investigate and isolate the sources of the overheads for the DNN models with different computational characteristics by studying various power modes on the device.
@@ 4.1 Hardware platform
Feature
AGX Orin
CPU Architecture
ARM Cortex A78AE
CPU ${}^{\dagger}$
12
CPU Frequency (${}^{\dagger}$
2200
GPU Architecture
Ampere
CUDA/Tensor Cores
2048/64
GPU Frequency (${}^{\dagger}$
1300
RAM (GB)
32, LPDDR5
Storage Interfaces
$\mu$SD, eMMC, NVMe, USB
Memory Frequency (${}^{\dagger}$
3200
Peak Power (W)
60
Price (USD)
\$$\$1999$
Form factor (mm)
$110\times 110\times 71.65$
${}^{\dagger}$ This is the maximum possible value across all power modes. Actual value depends on the power mode used
@@ 4.2 Software libraries
Model # Layers # Params ${}^{\dagger}$ Dataset # Samples Size on Disk Minibatch Size Used for
LeNet-5
7 [12]
$60k$
$4.4M$
MNIST
$60,000$
$46MB$
$16$
Train, Inf
MobileNet v3
20 [13]
$5.48M$
$225.4M$
GLD23k
$23,080$
$2.82GB$
$16$
Train, Inf
18 [14]
$11.68M$
$1.82G$
ImageNet
$50,000$
$6.74GB$
$16$
Train
ResNet-50
50 [14]
$26M$
$4G$
ImageNet
$50,000$
$6.74GB$
$16$
Inf
${}^{\dagger}$  As per the typical practice, FLOPS reported corresponds to a forward pass with minibatch size 1.  $*{}^{*}$ We use a smaller ResNet for training to avoid running out of memory
@@ 5 Results and Analysis
Figure 7: Compute time break-up for MobileNet training
@@ 5.1 DNN training
5.1.1 Lightweight training workloads see a significant increase in compute time when run in a container as opposed to bare metal
@@
In Fig 1a, we report average minibatch time as a stacked plot of fetch stall, compute time and logging overhead (minimal, not seen in plots). As seen from the plot, there is an increase in compute time from bare metal to container for all $3$ models. For LeNet, the compute time (2nd stack in the plot) increases from $3.91ms$ to $6.01ms$, an increase of $53.7\%$. MobileNet sees an increase from $82.12ms$ to $95.48ms$, which is $16.26\%$, and ResNet sees an increase from $50.53ms$ to $54.24ms$, which is $7.34\%$. In terms of FLOPS (Table 2), LeNet is the most lightweight of the $3$, followed by MobileNet and then ResNet. As can be seen, the increase in compute time follows the same pattern and is more for lightweight models LeNet and MobileNet.
5.1.2 This increase in compute time is due to CPU overheads of containerization
In order to localize the source of the compute increase, we look at violin plots of CPU, GPU utilization and power reported in Figs 2a,  2b and  2c respectively. Looking at MobileNet, we see that the median GPU utilization falls from $90\%$ on the bare metal to $77\%$ on the container, indicating that the GPU is spending time waiting. The lower GPU utilization also causes the median power for MobileNet to fall slightly from $43.7W$ to $40.8W$ as the GPU is major contributor to power. We also notice a higher median CPU utilization of $23\%$ on the container as opposed to $20.67\%$ on bare metal. This leads us to hypothesize that the CPU is the source of the compute overhead. For ResNet, the CPU utilization increase and GPU utilization decrease follow the same trend but are much less pronounced, indicating lesser compute overhead. Since LeNet is a very small model with very low GPU utilization, we do not observe a difference here. Most of the power for LeNet is contributedby the baseload as seen in Fig 2c, and we don’t see a difference here too.
5.1.3 This increase in compute time is further exacerbated when CPU frequency is lowered using a custom power mode, confirming that the CPU is the source of the overhead
To verify that the CPU is indeed the source of the overhead, we define $3$ custom power modes as listed in Table 3 and run the training workloads for these $3$ power modes apart from the default power mode (MAXN). Each of the power modes lowers one of CPU, GPU and memory frequencies while keeping the others the same as MAXN. LeNet and MobileNet container runs show a high sensitivity to lowering of CPU frequency more than GPU or memory–i.e. the compute time increases more sharply for container training when CPU frequency is lowered. In going from power mode MAXN to PM1_CPU, the compute time for MobileNet on bare metal increases from $82.12ms$ to $91ms$, an increase of $10.7\%$. In contrast, for the container, the increase in compute time is from $95.4ms$ to $130ms$, which is a much higher increase of $36.1\%$. We do not see this with other power modes. This proves conclusively that there is a CPU overhead of containerization which affects lightweight models more, and this effect gets worse if run in a power mode with reduced CPU frequency.
5.1.4 Why the overhead?
To further narrow down the reason for the CPU overhead, we traced the system calls for MobileNet training on bare metal and container using the Linux strace utility. We observed a larger number of cumulative system calls for the container. Since this pointed to a system call related overhead, we ran the container with seccomp turned off to observe the effect of Docker’s system call filtering. However, we did not observe any change in runtimes from the previous experiments. On a parallel track, we added more fine-grained profiling to the compute phase of the workload, recording times for forward pass, backward pass and the parameter update (optimizer.step in PyTorch). This breakup of the minibatch time is reported in Fig 7. We noticed that among the three compute steps, the parameter update using the SGD optimizer sees the most increase from $2.63ms$ to $17.63ms$, (a 6.7×6.7$\times$6.7 × increase as compared to $11.4\%$ increase for forward and $4.89\%$ increase for backward). We notice that container runs take longer after adding this fine-grained instrumentation and this might be due to asynchronous operations being affected by the torch.synchronize needed for recording the times. This needs further investiagtion. We also performed runs with 00 DataLoader workers to decouple the CPU overheads of the fetch/pre-process phase from the compute phase, but still observe the same overheads as before.
@@ 5.1.4 Why the overhead?
5.1.5 Energy consumption of the containerized training workload is higher than on bare metal, more so for lightweight models
@@
In Fig 1b, we report the energy consumed by the $3$ models for training as a stacked bar plot. The bottom stack of the bar represents the baseload, the energy it takes for the system to run with no workload. To measure this, we run our logging script for $10$min on bare metal with no workload running. Similarly, we spin up a container and run the logging script alone for $10$min. We scale both these energy values by the runtime of the actual workload on the bare metal and container respectively. The top stack reports the incremental energy spent for training i.e, energy over and above the baseload needed for model training. Again, we observe that there is an increase in total energy consumption for containerized training as compared to bare metal. This increase is around $48.89\%$ for LeNet, $11.11\%$ for MobileNet, and $6.84\%$ for ResNet, more significant for lightweight models.
5.1.6 This increase in energy for lightweight models is due to running for a longer time at a lower GPU utilization
As seen from Fig 1a, there is an increase in average minibatch time. The increase in energy is majorly due to the baseload increase of running for a longer time. For example, for MobileNet, we see a baseload increase of $30.76\%$ along with an incremental energy increase of $2.94\%$. This indicates that running for longer at a lower GPU utilization is not beneficial i.e, the power benefits obtained by running at a lower GPU utilization are outweighed by the energy consumed by running longer.
@@ 5.2 DNN inference
5.2.1 Lightweight inference workloads also see an increase in compute time, but the increase is less significant as compared to training
In Fig 3a, we report average minibatch time as a stacked bar plot of fetch stall, compute and overhead times. We observe that compute time increases by $30.52\%$ for LeNet and $12.93\%$ for MobileNet. ResNet does not have a noticeable increase in compute time (less than $1\%$). This can be explained by the fact that inference involves only the forward pass, and the parameter update which caused most of the overhead in training is not a part of inference workloads. We also perform experiments with the custom power modes for MobileNet and report the average minibatch times in Fig 6d. For the power mode with lowered CPU frequency (PM1_CPU), we observe that the compute time for the container increases by $33.72\%$ while for the bare metal it increases by $24.26\%$ as compared to MAXN. The compute increase for containerized inference is lesser compared to training.
5.2.2 Correspondingly, the energy increase is also modest for inference
In Fig 3b, we report the energy consumed by the $3$ models for inference as a stacked bar plot of baseload and incremental energy. The increase in total energy is $18.1\%$, $6.63\%$ and $2.68\%$ for LeNet, MobileNet and ResNet respectively, which is much lower than in training. This is due to the lower overheads of compute time in inference. Again, most of the energy increase comes from the baseload. For instance, MobileNet sees an increase of $20.03\%$ in the baseload and $0.5\%$ in incremental energy.
@@ 6 Conclusions and Future Work
In this paper, we closely examine and characterize containerized DNN training and inference workloads. We demonstrate that lightweight DNN models incur overheads in compute time and consequently energy when containerized. These overheads show up even more in power modes with lower CPU frequencies. This is relevant to federated learning workloads, which often involve lightweight models. In future, we plan to localize the overheads of containerization and also investigate the effect of containerization when inference and training are run concurrently, as this is an emerging usecase of continuous or lifelong learning.
@@ Acknowledgments
We thank members of the DREAM Lab, Tuhin Khare for plotting help and Ajay Nayak for inputs on Docker.

### v2

@@ Abstract
Edge devices have typically been used for DNN inferencing. The increase in the compute power of accelerated edges is leading to their use in DNN training also. As privacy becomes a concern on multi-tenant edge devices, Docker containers provide a lightweight virtualization mechanism to sandbox models. But their overheads for edge devices are not yet explored. In this work, we study the impact of containerized DNN inference and training workloads on an NVIDIA AGX Orin edge device and contrast it against bare metal execution on running time, CPU, GPU and memory utilization, and energy consumption. Our analysis shows that there are negligible containerization overheads for individually running DNN training and inference workloads.
@@ 1 Introduction
    We additionally investigate containerization performance for various power modes on the device.
@@
1.1 Updates from the HiPC Version based on Sensitivity to Software Versions
The results reported here are an updated version of those presented in our HiPC short paper [7]. We found that the PyTorch used by NVIDIA in L4T Containers for Jetson was based on a different commit of PyTorch 2.0.0 and not the release commit as indicated by the 2.0.0 tag (which was used in bare metal). This difference in Pytorch commits across bare metal and container caused a performance difference between bare metal and container. We have modified the Dockerfile and ensured that both bare metal and container use the same release commit of PyTorch 2.0.0, and the updated results show that there is no significant performance difference between bare metal and container for all the training and inference workloads that we study.
@@ 4.1 Hardware platform
Feature AGX Orin
CPU Architecture ARM Cortex A78AE
CPU Cores 12
CPU Frequency (MHz) 2200
GPU Architecture Ampere
CUDA/Tensor Cores 2048/64
GPU Frequency (MHz) 1300
RAM (GB) 32, LPDDR5
Storage Interfaces $\mu$SD, eMMC, NVMe, USB
Memory Frequency (MHz) 3200
Peak Power (W) 60
Price (USD) \$$\$1999$
Form factor (mm) $110\times 110\times 71.65$
 This is the maximum possible value across all power modes. Actual value depends on the power mode used
@@ 4.2 Software libraries
Model # Layers # Params FLOPS Dataset # Samples Size on Disk Minibatch Size Used for
LeNet-5 7 [13] $60k$ $4.4M$ MNIST $60,000$ $46MB$ $16$ Train, Inf
MobileNet v3 20 [14] $5.48M$ $225.4M$ GLD23k $23,080$ $2.82GB$ $16$ Train, Inf
ResNet-18 18 [15] $11.68M$ $1.82G$ ImageNet $50,000$ $6.74GB$ $16$ Train
ResNet-50 50 [15] $26M$ $4G$ ImageNet $50,000$ $6.74GB$ $16$ Inf
  As per the typical practice, FLOPS reported corresponds to a forward pass with minibatch size 1.   We use a smaller ResNet for training to avoid running out of memory
@@ 5.1 DNN training
5.1.1 Training workloads do not see an increase in compute time when run in a container as opposed to bare metal
@@
In Fig. 1a, we report average minibatch time as a stacked plot of fetch stall, compute time and logging overhead (minimal, not seen in plots). As seen from the plot, there is no increase in minibatch time from bare metal to container for all $3$ models. The fetch stall (1st stack) and compute times (2nd stack) also do not show any variation. For instance, for MobileNet, the average minibatch time is $88.89s$ on bare metal and $88.58s$ on container, a difference of $0.35\%$. The difference in minibatch time for ResNet is $0.58\%$. LeNet shows a difference of $2.5\%$, but this is because of small minibatch times and the absolute difference is only $0.14ms$.
5.1.2 CPU and GPU resource utilization and power consumption are also comparable across bare metal and container for training workloads
We look at violin plots of CPU, GPU utilization and power reported in Fig. 2a2b and 2c respectively. Looking at MobileNet in Fig 2a, we see that CPU utilization is $20.67\%$ on bare metal and $22.08\%$ on container. LeNet and ResNet follow a similar pattern, with the CPU utilization marginally (1-2%) higher for container. Similarly, from Fig 2b, we observe that GPU utilization for MobileNet are $91\%$ and $91.5\%$ for bare-metal and container respectively. ResNet and LeNet bare metal and container have identical GPU utilizations of $20\%$ and $99\%$ respectively. Considering Fig 2c, we see that median power for MobileNet is $43.21W$ on bare metal and $43.61W$ on container. LeNet has a median power of $18.1W$ and $18.5W$ on bare metal and container, and ResNet has a median power of $51.51W$ on both. Since the GPU is the major contributor to power, comparable utilizations across bare metal and container lead to the same power load.
5.1.3 Further, we do not see a significant increase in compute time even on custom power modes for training workloads.
We define $3$ custom power modes as listed in Table 3 and run the training workloads for these $3$ power modes apart from the default power mode (MAXN) to check if there are any overheads. Each of the power modes lowers one of CPU, GPU and memory frequencies while keeping the others the same as MAXN. For MobileNet, as seen from Fig. 6b, all power modes show similar minibatch times on bare metal and container. For ResNet, PM_MEM shows a very slight ($\approx 1\%$) increase in minibatch time on container. For LeNet, we do see some variation in minibatch times across bare metal and container. However, since the absolute minibatch times are very small (less than $10ms$) and show increased variability as indicated by the longer error bars, we do not consider this as a significant difference.
@@
5.1.4 Energy consumption of the containerized training workload matches bare metal.
@@
In Fig 1b, we report the energy consumed by the $3$ models for training as a stacked bar plot. The bottom stack of the bar represents the baseload, the energy it takes for the system to run with no workload. To measure this, we run our logging script for $10$min on bare metal with no workload running. Similarly, we spin up a container and run the logging script alone for $10$min. We scale both these energy values by the runtime of the actual workload on the bare metal and container respectively. The top stack reports the incremental energy spent for training i.e, energy over and above the baseload needed for model training.
We see some slight variations in the base and incremental energy components, but the difference in total energy is less than $1\%$ for all $3$ models. The total energy difference between container and bare metal is $0.58\%$, $0.75\%$ and $0.14\%$ for LeNet, MobileNet and ResNet respectively.
@@ 5.2 DNN inference
5.2.1 Inference workloads also do not see any increase in minibatch time
In Fig 3a, we report average minibatch time as a stacked bar plot of fetch stall, compute and overhead times. Similar to training, we observe that minibatch time is comparable across bare metal and container for inference workloads too. The difference in minibatch times across bare metal and container is $1.1\%$, $0.13\%$ and $0.04\%$ for LeNet, MobileNet and ResNet respectively.
We also perform experiments with the custom power modes for MobileNet and report the average minibatch times in Fig 6d.
The performance of bare metal and container is comparable across $3$ out of $4$ power modes. For the power mode with lowered CPU frequency (PM1_CPU), we observe that bare metal has a minibatch time of $37.65s$ and container $37.26$, a difference of $1\%$. However, we notice higher variability in minibatch times for this power mode as indicated by the longer error bars and this difference in time is within the error bar margin.
5.2.2 Correspondingly, inference energy is comparable across bare metal and container
In Fig 3b, we report the energy consumed by the $3$ models for inference as a stacked bar plot of baseload and incremental energy. We see slight variations in the base and incremental energy components, but the difference in total energy under $1\%$ for all $3$ models. The total energy in container is higher than bare metal by $0.81\%$, $0.86\%$ and $0.72\%$ for LeNet, MobileNet and ResNet respectively.
@@ 6 Conclusions and Future Work
In this paper, we closely examine and characterize containerized DNN training and inference workloads. We demonstrate that containerization overheads are minimal for both DNN training and inference workloads. Additionally, we also examine containerization overheads under different power modes and show that workload execution times do not vary. In future, we plan to investigate the effect of containerization when inference and training are run concurrently, as this is an emerging usecase of continuous or lifelong learning. We also plan to examine the overheads of Firecracker Micro VMs on edge devices as they offer enhanced security.
@@ Acknowledgments
We thank members of the DREAM Lab, particularly Beautlin S and Tuhin Khare for plotting help and Ajay Nayak from CSL Lab for inputs on Docker.